OpenStudy (anonymous):

The number of pages that Zev, Kelly, Marneisha, and Aleisha can read in a day is shown below. Zev read 15% of his 56-page book. Kelly read 12% of her 64-page book. Marneisha read 14% of his 68-page book Aleisha read 10% of her 72-page book. Who can read the greatest number of pages in a day?

OpenStudy (anonymous):

a - number of book's pages (72 for example) b - complete % (10% for example ) x - number of complete pages b= a/x So x = a/b Just calculate it

OpenStudy (anonymous):

WAIT I made a mistake b= x/a X= a*b

OpenStudy (anonymous):

Uhm, I am still kind of confused, could you walk me through the problem?